OpenStudy (anonymous):

integrate x-9/((x+5)(x-2))

OpenStudy (anonymous):

integrals are coming fast and furious. i think this one is partial fractions right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes definitely partial fractions because it is cooked to give nice integer answer. you want \[\frac{x-9}{(x+5)(x-2)}=\frac{a}{x+5}+\frac{b}{x-2}\]

OpenStudy (anonymous):

lets find a. if x = -5 the denominator of \[\frac{a}{x+5}\] is 0, so look at your original function, put your hand over the part that says \[x+5\] and replace x by -5 you get \[\frac{-5-9}{-5-2}=\frac{-14}{-7}=2\] so \[a=2\]

OpenStudy (anonymous):

likewise to find b, put your hand over the \[x-2\] part and replace x by 2 you get \[\frac{2-9}{2+5}=\frac{-7}{7}=-1\] so \[b=-1\]

OpenStudy (anonymous):

now your integral is \[\int\frac{2}{x+5}dx+\int\frac{-1}{x-2}dx\] which you do in your head and get \[2\ln(x+5)-\ln(x-2)\]

OpenStudy (anonymous):

all the work was writing \[\frac{x-9}{(x+5)(x-2)}=\frac{2}{x+5}+\frac{-1}{x-2}\]